# Eigenmatrix for unstructured sparse recovery

### 09 Feb 2024 v2

@@ 2.3. Real analytic case
To simplify the discussion, assume that $X$ is the interval $[-1,1]$, and we will comment on the general case later. Let us define for each $x$ the vector $\mathbf{g}(x)=[G(s_{j},x)]_{1\leq j\leq n_{s}}$. The first step is to construct $M$ such that $M\mathbf{g}(x)\approx x\mathbf{g}(x)$ for $x\in[-1,1]$.
Numerically, it is again more robust to use the normalized vector $\hat{\mathbf{g}}(x)=\mathbf{g}(x)/\|\mathbf{g}(x)\|$ and consider the modified condition
@@
We enforce this condition on a dense Chebyshev grid $\{a_{k}\}_{1\leq k\leq n_{a}}$ of size $n_{a}$ on the interval $[-1,1]$:
$M\hat{\mathbf{g}}(a_{k})\approx a_{k}\hat{\mathbf{g}}(a_{k}).$
Introduce the $n_{s}\times n_{s}$ matrix $W=[\hat{\mathbf{g}}(a_{k})]_{1\leq k\leq n_{a}}$ with columns $\hat{\mathbf{g}}(a_{k})$ as well as the $n_{a}\times n_{a}$ diagonal matrix $\Lambda=\mathrm{diag}(a_{k})$. The condition now reads
$MW\approx W\Lambda.$
This again suggests the following choice of the eigenmatrix for the real analytic case
$M:=W\Lambda W^{\dagger},$
where the pseudoinverse $W^{\dagger}$ is computed by thresholding the singular values of $W$ below $\sigma$.
With $M$ available, we define the vector $\tilde{\mathbf{u}}:=[\tilde{u}_{j}]_{1\leq j\leq n_{s}}$, form the system
$\begin{bmatrix}\tilde{\mathbf{u}}&M\tilde{\mathbf{u}}&\ldots&M^{d}\tilde{\mathbf{u}}\end{bmatrix}\tilde{\mathbf{p}}=0\quad\text{with}\quad\tilde{\mathbf{p}}=\begin{bmatrix}\tilde{p}_{0}\\ \vdots\\ \tilde{p}_{d}\end{bmatrix},$
and solve for $\tilde{\mathbf{p}}$ as the null space vector. The roots of the polynomial $\tilde{p}(x):=\tilde{p}_{0}+\tilde{p}_{1}x+\cdots+\tilde{p}_{d}x^{d}$ are the estimators $\{\tilde{x}_{k}\}$ for the spike locations and a least square solve gives $\{\tilde{w}_{k}\}$ as the estimators for the spike weights.
@@
We claim that, for real analytic kernels $G(s,x)$, enforcing the condition at the Chebyshev grid $\{a_{k}\}$ is sufficient. To see this,
$M\mathbf{g}(x)\approx M\left(\sum_{k}c_{k}(x)\mathbf{g}(a_{k})\right)=\sum_{k}c_{k}(x)M\mathbf{g}(a_{k})\approx\sum_{k}c_{k}(x)(a_{k}\mathbf{g}(a_{k}))\approx x\mathbf{g}(x),$
where $c_{k}(x)$ is the Chebyshev quadrature for $x$ associated with grid $\{a_{k}\}$. Here, the first and third approximations use the convergence property of the Chebyshev quadrature for analytic functions $\mathbf{g}(x)$ and $x\mathbf{g}(x)$, and the second approximation directly comes from $M\mathbf{g}(a_{k})\approx a_{k}\mathbf{g}(a_{k})$.
@@ 3. Numerical results
This section applies the eigenmatrix approach to the unstructured sparse recovery problems mentioned in Section 1. In all examples, the size of the uniform or Chebyshev grid is $n_{a}=1024$. The spike weights $\{w_{k}\}$ are set to be $1$ and the noises $\{Z_{j}\}$ are Gaussian.
@@
Figure 1. Rational approximation. $G(s,x)=\frac{1}{s-x}$. $X=\mathbb{D}_{1}$. $\{s_{j}\}$ are random points outside the unit disk, each with a modulus between $1.2$ and $2.2$. $n_{s}=40$. In the three columns, $\sigma$ equals to $10^{-2}$, $10^{-3}$, and $10^{-4}$, respectively. Blue, green, and red stand for the exact solution, the result before postprocessing, and the one after postprocessing. $n_{x}=4$. Top row: the test with well-separated locations. Bottom row: the test with two nearby locations.
Figure 1 summarizes the experimental results. The three columns correspond to noise levels $\sigma$ equal to $10^{-2}$, $10^{-3}$, and $10^{-4}$. In each plot, the blue markers are the exact solution $\{x_{k}\}$, the green ones are the solution of the eigenmatrix approach before the postprocessing, and the red ones are the solution after the postprocessing.
Two tests are performed. In the first one (top row), $n_{x}=4$ and $\{x_{k}\}$ are well-separated from each other. The plots show accurate recovery of the spike locations from all $\sigma$ values. In the second one (bottom row), $n_{x}=4$ and two of the spike locations are close to each other. In this harder case, the reconstruction at $\sigma=10^{-2}$ shows a noticeable error, while the results for $\sigma=10^{-3}$ and $\sigma=10^{-4}$ are accurate. The plots also suggest that the eigenmatrix approach results in fairly accurate (green) initial guesses for the postprocessing step.
@@
Figure 2. Spectral function approximation. $G(s,x)=\frac{1}{s-x}$. $X=[-1,1]$. $\{s_{j}\}$ is the Matsubara grid from $-\frac{(2N-1)\pi}{\beta}i$ to $\frac{(2N-1)\pi}{\beta}i$ with $\beta=100$ and $N=128$. In the three columns, $\sigma$ equals to $10^{-2}$, $10^{-3}$, and $10^{-4}$, respectively. Blue, green, and red stand for the exact solution, the result before postprocessing, and the one after postprocessing. $n_{x}=4$. Top row: the test with well-separated locations. Bottom row: the test with two nearby locations.
Figure 2 summarizes the experimental results. The three columns correspond to $\sigma$ equal to $10^{-2}$, $10^{-3}$, and $10^{-4}$, respectively. The blue, green, and red spikes denote the exact solution of the reconstructions before and after postprocessing, respectively.
Two tests are performed. In the first one (top row), $n_{x}=4$ and $\{x_{k}\}$ are well-separated. The reconstructions are accurate for all $\sigma$ values. In the second one (bottom row), $n_{x}=4$ and two of the spike locations are within $0.1$ distance from each other. In this harder case, the reconstructions remain accurate for all $\sigma$ values. Notice that the eigenmatrix provides a sufficiently accurate initial guess for postprocessing.
@@
Figure 3. Fourier inversion. $G(s,x)=\exp(\pi isx)$. $X=[-1,1]$. $\{s_{j}\}$ are randomly chosen points in $[-5,5]$. $n_{s}=128$. In the three columns, $\sigma$ equals to $10^{-2}$, $10^{-3}$, and $10^{-4}$, respectively. Blue, green, and red stand for the exact solution, the result before postprocessing, and the one after postprocessing. $n_{x}=4$. Top row: the test with well-separated locations. Bottom row: the test with two nearby locations.
Figure 3 summarizes the experimental results. The three columns correspond to $\sigma$ equal to $10^{-2}$, $10^{-3}$, and $10^{-4}$. The blue, green, and red spikes denote the exact solution of the reconstructions before and after postprocessing, respectively.
Two tests are performed. In the first one (top row), $n_{x}=4$ and $\{x_{k}\}$ are well-separated. The reconstructions are accurate for all $\sigma$ values. In the second one (bottom row), $n_{x}=4$ and two of the spike locations are within $0.1$ distance from each other. The reconstructions are also accurate for all $\sigma$ values. The eigenmatrix is again able to provide sufficient accurate initial guesses for the postprocessing step.
@@
Figure 4. Laplace inversion. $G(s,x)=\exp(-sx)$. $X=[0,2]$. $\{s_{j}\}$ are random samples in $[0,10]$. $n_{s}=100$. In the three columns, $\sigma$ equals to $10^{-5}$, $10^{-6}$, and $10^{-7}$, respectively. Blue, green, and red stand for the exact solution, the result before postprocessing, and the one after postprocessing. $n_{x}=4$. Top row: the test with well-separated locations. Bottom row: the test with two nearby locations.
Figure 4 summarizes the experimental results. The inverse Laplace transform is well-known for its sensitivity to noise. As a result, significantly smaller noise magnitudes are used in this example: the three columns correspond to $\sigma$ equal to $10^{-5}$, $10^{-6}$, and $10^{-7}$. The blue, green, and red spikes still denote the exact solution of the reconstructions before and after postprocessing, respectively.
Two tests are performed. In the first one (top row), $n_{x}=4$ and $\{x_{k}\}$ are well-separated. The reconstructions are acceptable for $\sigma=10^{-6}$ and accurate for $\sigma=10^{-7}$. In the second harder test (bottom row), $n_{x}=4$ and two of the spike locations are within $0.25$ distance from each other. The reconstructions provide reasonable reconstructions at $\sigma=10^{-7}$, but significant errors for larger $\sigma$ values.
@@
Figure 5. Sparse deconvolution. $G(s,x)=\frac{1}{1+\gamma(s-x)^{2}}$ with $\gamma=4$. $X=[-1,1]$. $\{s_{j}\}$ are random samples from $[-5,5]$. $n_{s}=100$. In the three columns, $\sigma$ equals to $10^{-2}$, $10^{-3}$, and $10^{-4}$, respectively. Blue, green, and red stand for the exact solution, the result before postprocessing, and the one after postprocessing. $n_{x}=4$. Top row: the test with well-separated locations. Bottom row: the test with two nearby locations.
Figure 5 summarizes the experimental results. The three columns correspond to $\sigma$ equal to $10^{-2}$, $10^{-3}$, and $10^{-4}$. The blue, green, and red spikes denote the exact solution of the reconstructions before and after postprocessing, respectively.
Two tests are performed. In the first one (top row), $n_{x}=4$ and $\{x_{k}\}$ are well-separated. The reconstructions are reasonable for $\sigma=10^{-2}$ and accurate for the smaller $\sigma$ values. In the second one (bottom row), $n_{x}=4$ and two of the spike locations are within $0.1$ distance from each other. The reconstructions are accurate for $\sigma$ equal to $10^{-3}$ and $10^{-4}$.
@@
The numerical experience suggests two lessons important for accurate reconstruction. First, it is important to fully exploit the prior information about the support of the spikes, i.e., making the candidate parameter set $X$ as compact as possible. Second, using the Chebyshev and uniform grids ensures that $M\mathbf{g}(x)\approx x\mathbf{g}(x)$ up to a high accuracy numerically.
@@ 4. Discussions
This paper introduces the eigenmatrix construction for unstructured sparse recovery problems. As a data-driven approach, it assumes no structure on the sample locations and offers a rather unified framework for such sparse recovery problems. There are several directions for future work.
@@
  • Once the eigenmatrix is constructed, the recovery algorithm presented above follows the classical Prony’s method. An immediate extension is to apply the eigenmatrix with other Prony-type algorithms, such as MUSIC, ESPRIT, and the matrix pencil method.

### 08 Mar 2024 v4

@@ 3.3. Real analytic case
To simplify the discussion, assume that $X$ is the interval $[-1,1]$, and we will comment on the general case later. Let us define for each $x$ the vector $\mathbf{g}(x)=[G(s_{j},x)]_{1\leq j\leq n_{s}}$. The first step is to construct $M$ such that $M\mathbf{g}(x)\approx x\mathbf{g}(x)$ for $x\in[-1,1]$. Numerically, it is again more robust to use the normalized vector $\hat{\mathbf{g}}(x)=\mathbf{g}(x)/\|\mathbf{g}(x)\|$ and consider the modified condition
@@
We enforce this condition on a Chebyshev grid $\{a_{t}\}_{1\leq t\leq n_{a}}$ of size $n_{a}$ on the interval $[-1,1]$:
$M\hat{\mathbf{g}}(a_{t})\approx a_{t}\hat{\mathbf{g}}(a_{t}).$
Introduce the $n_{s}\times n_{s}$ matrix $\hat{G}=[\hat{\mathbf{g}}(a_{t})]_{1\leq t\leq n_{a}}$ with columns $\hat{\mathbf{g}}(a_{t})$ as well as the $n_{a}\times n_{a}$ diagonal matrix $\Lambda=\mathrm{diag}(a_{t})$. The condition now reads
$M\hat{G}\approx\hat{G}\Lambda.$
When the columns of $\hat{G}$ are numerically linearly independent, this again suggests the following choice of the eigenmatrix for the real analytic case
$M:=\hat{G}\Lambda\hat{G}^{\dagger},$
where the pseudoinverse $\hat{G}^{\dagger}$ is computed by thresholding the singular values of $\hat{G}$.
@@
We claim that, for real analytic kernels $G(s,x)$, enforcing the condition at the Chebyshev grid $\{a_{t}\}$ is sufficient. To see this,
$M\mathbf{g}(x)\approx M\left(\sum_{k}c_{t}(x)\mathbf{g}(a_{t})\right)=\sum_{t}c_{t}(x)M\mathbf{g}(a_{t})\approx\sum_{t}c_{t}(x)(a_{t}\mathbf{g}(a_{t}))\approx x\mathbf{g}(x),$
where $c_{t}(x)$ is the Chebyshev quadrature for $x$ associated with grid $\{a_{t}\}$. Here, the first and third approximations use the convergence property of the Chebyshev quadrature for analytic functions $\mathbf{g}(x)$ and $x\mathbf{g}(x)$, and the second approximation directly comes from $M\mathbf{g}(a_{t})\approx a_{t}\mathbf{g}(a_{t})$.
@@
3.4. Putting together
With the eigenmatrix $M$ available, the rest is similar to Prony’s method and the ESPRIT algorithm. Define the vector $\tilde{\mathbf{u}}:=[\tilde{u}_{j}]_{1\leq j\leq n_{s}}$ from the noisy sample values.
For Prony’s method, consider
$\begin{bmatrix}\tilde{\mathbf{u}}&M\tilde{\mathbf{u}}&\ldots&M^{n_{x}}\tilde{\mathbf{u}}\end{bmatrix}\approx\begin{bmatrix}\mathbf{g}(x_{1})&\ldots&\mathbf{g}(x_{n_{x}})\end{bmatrix}\begin{bmatrix}w_{1}&&\\ &\ddots&\\ &&w_{n_{x}}\end{bmatrix}\begin{bmatrix}1&x_{1}&\ldots&x_{1}^{n_{x}}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n_{x}}&\ldots&x_{n_{x}}^{n_{x}}\end{bmatrix}.$
Let $\tilde{\mathbf{p}}$ be a non-zero vector in its null-space
$\begin{bmatrix}\tilde{\mathbf{u}}&M\tilde{\mathbf{u}}&\ldots&M^{n_{x}}\tilde{\mathbf{u}}\end{bmatrix}\tilde{\mathbf{p}}=0\quad\text{with}\quad\tilde{\mathbf{p}}=\begin{bmatrix}\tilde{p}_{0}\\ \vdots\\ \tilde{p}_{n_{x}}\end{bmatrix}$
Therefore,
$\begin{bmatrix}1&x_{1}&\ldots&x_{1}^{n_{x}}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n_{x}}&\ldots&x_{n_{x}}^{n_{x}}\end{bmatrix}\begin{bmatrix}\tilde{p}_{0}\\ \vdots\\ \tilde{p}_{n_{x}}\end{bmatrix}\approx 0.$
The roots of the polynomial $\tilde{p}(x):=\tilde{p}_{0}+\tilde{p}_{1}x+\cdots+\tilde{p}_{n_{x}}x^{n_{x}}$ provide the estimators $\{\tilde{x}_{k}\}$ for $\{x_{k}\}$.
For the ESPRIT algorithm, consider the matrix
$\begin{bmatrix}\tilde{\mathbf{u}}&M\tilde{\mathbf{u}}&\ldots&M^{\ell}\tilde{\mathbf{u}}\end{bmatrix}\approx\begin{bmatrix}\mathbf{g}(x_{1})&\ldots&\mathbf{g}(x_{n_{x}})\end{bmatrix}\begin{bmatrix}w_{1}&&\\ &\ddots&\\ &&w_{n_{x}}\end{bmatrix}\begin{bmatrix}1&x_{1}&\ldots&x_{1}^{\ell}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n_{x}}&\ldots&x_{n_{x}}^{\ell}\end{bmatrix}$
with $\ell>n_{x}$. Let $\tilde{U}\tilde{S}\tilde{V}^{*}$ be the rank-$n_{x}$ truncated SVD of this matrix. The matrix $\tilde{V}^{*}$ satisfies
$\tilde{V}^{*}\approx P\begin{bmatrix}1&x_{1}&\ldots&x_{1}^{\ell}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n_{x}}&\ldots&x_{n_{x}}^{\ell}\end{bmatrix}$
where $P$ is an unknown non-degenerate $n_{x}\times n_{x}$ matrix. Let $\tilde{Z}_{0}$ and $\tilde{Z}_{1}$ be the submatrices of $\tilde{Z}$ by excluding the first column and the last column, respectively, i.e.,
$\tilde{Z}_{0}\approx P\begin{bmatrix}1&\ldots&x_{1}^{\ell-1}\\ \vdots&\ddots&\vdots\\ 1&\ldots&x_{n_{x}}^{\ell-1}\end{bmatrix},\quad\tilde{Z}_{1}\approx P\begin{bmatrix}x_{1}&\ldots&x_{1}^{\ell}\\ \vdots&\ddots&\vdots\\ x_{n_{x}}&\ldots&x_{n_{x}}^{\ell}\end{bmatrix}.$
By introducing
$\tilde{Z}_{1}(\tilde{Z}_{0})^{+}\approx P\begin{bmatrix}x_{1}&&\\ &\ddots&\\ &&x_{n_{x}}\end{bmatrix}P^{-1},$
one can get estimates $\{\tilde{x}_{k}\}$ for $\{x_{k}\}$ by computing the eigenvalues of $\tilde{Z}_{1}(\tilde{Z}_{0})^{+}$.
With $\{\tilde{x}_{k}\}$ available, the least square solve
gives the estimators $\{\tilde{w}_{k}\}$ for $\{w_{k}\}$ for both methods.
@@ 4. Numerical results
This section applies the eigenmatrix approach to the unstructured sparse recovery problems mentioned in Section 1. In all examples, the spike weights $\{w_{k}\}$ are set to be $1$ and the noises $\{Z_{j}\}$ are Gaussian.
Once the eigenmatrix $M$ is constructed, the reported numerical results are obtained using ESPRIT. The results of Prony’s method are similar but slightly less robust. In each plot, blue, green, and red stand for the exact solution, the result before postprocessing, and the one after postprocessing.
@@
Figure 2. Rational approximation. $G(s,x)=\frac{1}{s-x}$. $X=\mathbb{D}$. $\{s_{j}\}$ are random points outside the unit disk, each with a modulus between $1.2$ and $2.2$. $n_{s}=40$. Columns: $\sigma$ equals to $10^{-2}$, $10^{-3}$, and $10^{-4}$. Rows: the easy test (well-separated) and the hard test (with two nearby spikes).
Figure 2 summarizes the experimental results. $n_{a}=32$. The three columns correspond to noise levels $\sigma$ equal to $10^{-2}$, $10^{-3}$, and $10^{-4}$. Two tests are performed. In the first one (top row), $\{x_{k}\}$ are well-separated from each other. The plots show accurate recovery of the spike locations from all $\sigma$ values. In the second one (bottom row), two of the spike locations are close to each other: the reconstruction at $\sigma=10^{-2}$ shows a noticeable error, while the results for $\sigma=10^{-3}$ and $\sigma=10^{-4}$ are accurate. The plots also suggest that the eigenmatrix approach results in fairly accurate (green) initial guesses for the postprocessing step.
@@
Figure 3. Spectral function approximation. $G(s,x)=\frac{1}{s-x}$. $X=[-1,1]$. $\{s_{j}\}$ is the Matsubara grid from $-\frac{(2N-1)\pi}{\beta}i$ to $\frac{(2N-1)\pi}{\beta}i$ with $\beta=100$ and $N=128$. Columns: $\sigma$ equals to $10^{-2}$, $10^{-3}$, and $10^{-4}$. Rows: the easy test (well-separated) and the hard test (with two nearby spikes).
Figure 3 summarizes the experimental results. $n_{a}=32$. The three columns correspond to $\sigma$ equal to $10^{-2}$, $10^{-3}$, and $10^{-4}$, respectively. Two tests are performed. In the first one (top row), $\{x_{k}\}$ are well-separated. The reconstructions are accurate for all $\sigma$ values. In the second one (bottom row), two of the spike locations are within $0.1$ distance from each other. In this harder case, the reconstructions also remain accurate for all $\sigma$ values. Notice that the eigenmatrix provides a sufficiently accurate initial guess for postprocessing.
@@
Figure 4. Fourier inversion. $G(s,x)=\exp(\pi isx)$. $X=[-1,1]$. $\{s_{j}\}$ are randomly chosen points in $[-5,5]$. $n_{s}=128$. Columns: $\sigma$ equals to $10^{-2}$, $10^{-3}$, and $10^{-4}$, respectively. Rows: the easy test (well-separated) and the hard test (with two nearby spikes).
Figure 4 summarizes the experimental results. $n_{a}=32$. The three columns correspond to $\sigma$ equal to $10^{-2}$, $10^{-3}$, and $10^{-4}$. Two tests are performed. In the first one (top row), $\{x_{k}\}$ are well-separated. The reconstructions are accurate for all $\sigma$ values. In the second one (bottom row), two of the spike locations are within $0.1$ distance from each other. The reconstructions are also accurate for all $\sigma$ values. The eigenmatrix is again able to provide sufficient accurate initial guesses for the postprocessing step.
@@
Figure 5. Laplace inversion. $G(s,x)=\exp(-sx)$. $X=[0.1,2.1]$. $\{s_{j}\}$ are random samples in $[0,10]$. $n_{s}=100$. Columns: $\sigma$ equals to $10^{-5}$, $10^{-6}$, and $10^{-7}$, respectively. Rows: the easy test (well-separated) and the hard test (with two nearby spikes).
Figure 5 summarizes the experimental results. $n_{a}=32$. The inverse Laplace transform is well-known for its sensitivity to noise. As a result, significantly smaller noise magnitudes are used in this example: the three columns correspond to $\sigma$ equal to $10^{-5}$, $10^{-6}$, and $10^{-7}$. Two tests are performed. In the first one (top row), $\{x_{k}\}$ are well-separated. The reconstructions are acceptable for $\sigma=10^{-6}$ and accurate for $\sigma=10^{-7}$. In the second harder test (bottom row), two of the spike locations are within $0.25$ distance from each other. The reconstructions provide reasonable reconstructions at $\sigma=10^{-7}$, but significant errors for larger $\sigma$ values.
@@
Figure 6. Sparse deconvolution. $G(s,x)=\frac{1}{1+\gamma(s-x)^{2}}$ with $\gamma=4$. $X=[-1,1]$. $\{s_{j}\}$ are random samples from $[-5,5]$. $n_{s}=100$. Columns: $\sigma$ equals to $10^{-2}$, $10^{-3}$, and $10^{-4}$, respectively. Rows: the easy test (well-separated) and the hard test (with two nearby spikes).
Figure 6 summarizes the experimental results. $n_{a}=32$. The three columns correspond to $\sigma$ equal to $10^{-2}$, $10^{-3}$, and $10^{-4}$. Two tests are performed. In the first one (top row), $\{x_{k}\}$ are well-separated. The reconstructions are reasonable for $\sigma=10^{-2}$ and accurate for the smaller $\sigma$ values. In the second one (bottom row), two of the spike locations are within $0.1$ distance from each other. The reconstructions are accurate for $\sigma$ equal to $10^{-3}$ and $10^{-4}$.
@@
The numerical experience suggests two lessons important for accurate reconstruction. First, it is important to fully exploit the prior information about the support of the spikes, i.e., making the candidate parameter set $X$ as compact as possible. Second, using the Chebyshev grid (for the real case) and the uniform grid (for the complex case) ensures that $M\mathbf{g}(x)\approx x\mathbf{g}(x)$ up to a high accuracy numerically.
@@ 5. Discussions
This note introduces the eigenmatrix construction for unstructured sparse recovery problems. It assumes no structure on the sample locations and offers a rather unified framework for such sparse recovery problems. This note is only an exploratory study of the data-driven approach for unstructured sparse recovery, and there are several clear directions for future work.
  • Providing a more principled way of choosing the size of the grid $\{a_{t}\}$ and the thresholding value for computing $M$.
@@
  • Once the eigenmatrix is constructed, the recovery algorithm presented above follows Prony’s method and the ESPRIT algorithm. An immediate extension is to combine the eigenmatrix with other algorithms, such as MUSIC and the matrix pencil method.